# fastrerandomize: An R Package for Fast Rerandomization Using Accelerated Computing

### v1

@@ Abstract
The fastrerandomize R package provides hardware-accelerated tools for performing rerandomization and randomization testing in experimental research. Using a JAX backend, the package enables exact rerandomization inference even for large experiments with hundreds of billions of possible randomizations. Key functionalities include generating pools of acceptable rerandomizations based on covariate balance, conducting exact randomization tests, and performing pre-analysis evaluations to determine optimal rerandomization acceptance thresholds. Through batched processing and GPU acceleration, fastrerandomize achieves substantial performance gains compared to existing implementations, making previously intractable designs computationally feasible. The package therefore extends the randomization-based inference toolkit in R, allowing researchers to efficiently implement more stringent rerandomization designs and conduct valid inference even with large sample sizes or in high-dimensional settings.
@@
Rerandomization, randomization tests, experimental design, hardware acceleration, R
@@ Introduction
Rerandomization is a powerful strategy for ensuring strong covariate balance in experimental research. By repeatedly drawing treatment assignments until a prespecified balance criterion is met, rerandomization helps improve the precision of estimated treatment effects relative to pure random assignment (Morgan and Rubin, 2012; Li et al., 2018). Despite its advantages, rerandomization poses a formidable computational challenge: as the size of an experiment grows, the space of possible assignments can quickly reach hundreds of billions, complicating both the generation of acceptable randomizations and the conduct of valid inference. Although theoretical work has established the statistical validity of rerandomization, user-friendly software that efficiently implements these methods in large-scale settings has, until now, been more limited.
The fastrerandomize package addresses this need by providing hardware-accelerated tools for rerandomization and randomization testing within R. By leveraging JAX, fastrerandomize supports hardware backends such as CPU, CUDA, and METAL, enabling researchers to conduct exact randomization-based inference even at large experimental scales. Key functionalities include (1) generating pools of acceptable assignments based on desired balance thresholds, (2) running exact randomization tests with the accepted pools, and (3) performing pre-analysis evaluations to guide the choice of optimal rerandomization thresholds.
@@
In what follows, we provide an overview of the statistical framework underpinning fastrerandomize, detail its core functionalities, and illustrate how it may be applied in practice. After literature review in Section 1, Section 2 outlines the primary methodologies implemented in the package, Section 3 describes its software architecture and usage. Sections 4 and 5 offer examples demonstrating both basic and more advanced analyses. We conclude with potential directions for future development in Section 7.
@@
Figure 1: Core workflow of the fastrerandomize package for conducting rerandomization and randomization testing. The process begins with covariate data and proceeds through five main steps: (1) generating candidate randomizations using compiled code, (2) checking covariate balance with GPU-accelerated computations, (3) accepting or rejecting assignments using Monte Carlo methods for large experiments, (4) conducting exact randomization tests using the pool of accepted randomizations, and (5) optionally computing fiducial intervals. Steps are implemented with hardware acceleration to enable efficient processing of large-scale experiments.
@@ 1 Literature Review and Related Packages
Over the past decade, an expanding body of open-source software has emerged to facilitate experimental design, with a particular focus on random assignment procedures and randomization-based inference. Many of these tools incorporate rerandomization techniques to improve covariate balance across treatment groups, thereby enhancing the statistical power of a study (Morgan and Rubin, 2012; Li et al., 2018). Despite the proliferation of solutions, existing packages often address only narrow classes of experiments and frequently encounter computational bottlenecks in large-scale settings with many units or covariates. This section surveys some of the most widely used packages for general random assignment, randomization testing, and rerandomization, before discussing the core challenges that persist in these domains.
@@ 1.1 General Random Assignment and Testing Packages
A variety of open-source packages are available for general-purpose random assignment and randomization-based testing. The randomizr package, for instance, streamlines the assignment of study units to diverse treatment groups (Coppock, 2024). Another solution, RATest, incorporates randomization tests applicable to multiple experimental design scenarios and offers sample datasets that illustrate, for example, permutation testing in regression discontinuity contexts (Olivares-Gonzalez and Sarmiento-Barbieri, 2017).
@@ 1.3 Challenges and Gaps
Although these packages advance the design and analysis of experiments, several limitations constrain their application to large-scale studies. One persistent concern is that many solutions can verify or test an existing randomization but provide limited functions for the initial task of constructing assignment schemes automatically. This requires researchers to generate and manage their own candidate pools of randomizations.
Scalability also presents a bottleneck in rerandomization: as experiments grow to include more units and covariates, generating, evaluating, and storing large pools of acceptable randomizations becomes computationally intensive. Current software packages often struggle when handling scenarios where investigators seek stringent acceptance thresholds with hundreds of billions of potential randomizations, or in experiments involving hundreds or thousands of covariates—a scenario becoming increasingly common with the rise of causal machine learning involving high-dimensional text (Keith et al., 2020), image (Jerzak et al., 2023), or network data (Ogburn et al., 2024).
While modern hardware offers parallelism gains through accelerated computing, existing implementations have not leveraged these advances. The fastrerandomize package addresses these limitations by automating balanced assignment generation and harnessing this acceleration, thereby making rerandomization and randomization-based inference both computationally efficient and practically feasible for large-scale experimental research.
@@ 2 Statistical Methods
Conversely, overly lenient thresholds can diminish the advantages of rerandomization by accepting many poorly balanced designs, causing the method to converge back to classical complete randomization. Accordingly, researchers often iterate over plausible acceptance probabilities in “pre-analysis” simulations to gauge how different thresholds affect downstream power, $p$-values, or interval widths (Branson et al., 2024; Kapelner et al., 2022). Such simulations also help ensure that the final pool of accepted randomizations is large enough for reliable test-based inferences and for constructing credible fiducial intervals, particularly in high-dimensional covariate settings (e.g., text, image, or network data (Keith et al., 2020; Jerzak et al., 2023; Ogburn et al., 2024)). [^1] In high-dimensional applications, one can improve numerical stability by incorporating penalized or regularized covariance estimates (e.g., diagonal or ridge-regularized inverses) into the balance statistic, ensuring the acceptance criterion continues to reflect meaningful imbalances.
@@ 3 Computational Methods
Implementing rerandomization at scale involves navigating an inherently combinatorial process, in which the total number of possible treatment assignments can be enormous. To address these computational challenges, fastrerandomize employs a hybrid of exact enumeration (for smaller designs) and Monte Carlo sampling (for large designs) while leveraging batched processing for memory efficiency. Central to this approach is the use of key-based storage, which replaces the explicit storage of every randomization vector with small keys that allow the randomizations to be regenerated on demand. Additionally, approximate techniques (e.g., diagonal covariance approximations) expedite balance checking in high-dimensional covariate spaces. Finally, the package exploits just-in-time (JIT) compilation and hardware acceleration via GPUs or specialized backends to further reduce run times. The methods outlined here enable fastrerandomize to handle rerandomization tasks that would be computationally prohibitive using naive approaches, even when the number of units, covariates, or candidate randomizations is very large.
@@ 3.1 Exact Enumeration and Batched Processing
The fastrerandomize package addresses a central challenge in rerandomization: balancing exactness with practical computational demands. To tackle this challenge, the package provides two primary strategies for generating and storing randomizations—exact enumeration and Monte Carlo sampling—each designed to accommodate different experimental sizes and memory constraints. Both approaches are implemented in JAX and exploit a batched workflow to reduce memory overhead, while R-facing functions manage data validation, type conversion, and user interaction. This integrated approach enables researchers to efficiently construct pools of acceptable randomizations across a range of experimental scales.
@@
For smaller studies where the total number of possible assignments, $\binom{n}{n_{\textrm{Treated}}},$ is sufficiently small (e.g., up to a few hundred million), it is possible to exhaustively enumerate every distinct randomization. Users can invoke exact enumeration by specifying:
generate_randomizations(
n_units = 10,
n_treated = 5,
X = matrix(rnorm(10*3), nrow=10),
randomization_accept_prob = 0.1,
randomization_type = "exact"
This approach systematically iterates over all assignments, providing the maximum level of precision and guaranteeing that the entire randomization space is captured. Although exact enumeration is attractive from a statistical perspective, it becomes infeasible for even moderate sample sizes (e.g., $n=50$), because the number of possible assignments grows combinatorially.
@@
For larger experiments, fastrerandomize supports a Monte Carlo mechanism that draws candidate assignments in batches and evaluates their covariate balance:
big_covariate_matrix <- matrix(rnorm(500*100), nrow=500)
generate_randomizations(
n_units = nrow(big_covariate_matrix),
n_treated = round(nrow(big_covariate_matrix)/2),
X = big_covariate_matrix,
randomization_type = "monte_carlo",
max_draws = 5e5,
randomization_accept_prob = 0.001
Under Monte Carlo sampling, candidate assignments are randomly generated and checked against a user-specified criterion (e.g., Hotelling’s $T^{2}$). Those randomizations meeting the threshold are accepted. This approach can rapidly explore massive assignment spaces without enumerating them in full and is typically favored for large $n$.
Batched Processing and Memory Management.
Regardless of whether the randomizations are generated by exact enumeration or Monte Carlo sampling, fastrerandomize relies on batching to handle large-scale experiments efficiently:
  • Chunked Generation. Both the exhaustive listing of assignments and the Monte Carlo draws can be subdivided into user-defined batches (controlled by batch_size). This prevents the software from loading all candidate randomizations at once into memory.
  • Flexible Return Options. Users can further reduce memory loads by writing accepted assignments directly to disk rather than returning them to R (using file in generate_randomizations()). This is especially helpful when generating hundreds of thousands or millions of accepted randomizations.
By selecting exact enumeration for small designs and Monte Carlo with batching for large ones, analysts can scale rerandomization from modest experiments to those involving thousands of units. This unified approach provides both the precision afforded by thorough enumeration and the computational tractability of sample-based generation, all while leveraging batched execution to prevent overwhelming system memory.
Moreover, in large-scale experiments involving Monte Carlo sampling, storing all candidate randomizations in memory quickly becomes infeasible as $n$ and the total number of candidate randomizations grow. To address this, our implementation employs a streamlined approach where each batch of candidate randomizations is generated and evaluated, with only key information being preserved for the final evaluation step. This keeps the memory footprint low. The result is a scalable procedure that supports large experiments, since we avoid loading or caching an unwieldy number of candidate assignments in a single data structure. This key-generation process is visualized in Figure 3.
@@
A Brief Model of Memory Gains from Key-Only Storage
@@
Instead of storing every candidate randomization vector directly, one may store only an associated key $\kappa^{(m)}$ for each $W^{(m)}$. Here, $\kappa^{(m)}\in\{\mathbb{N}_{0}\}^{L}$ is a short seed (or split key) that can be used by a pseudorandom number generator (PRNG) to regenerate the corresponding vector $W^{(m)}$ on demand:
@@
For an experiment with 1000 units, this ratio would be $1000/2$, indicating memory usage improvements of 500 times. At any later point, to reconstruct the accepted $W^{(m)}$ values for final analysis, we simply apply the PRNG again to these stored keys.
@@
3.2 Approximations
When evaluating multivariate covariate balance, the package provides a default threshold function based on Hotelling’s $T^{2}$. For very large matrices, an approximate inverse of the sample covariance matrix can be used (approximate_inv = TRUE), reducing computational overhead from matrix inversion while still enabling balance checking.
3.3 Just-in-Time Compilation with Precision Management
Each internal function (such as FastHotel2T2 and FastDiffInMeans) is decorated with JAX just-in-time (JIT) compilation, which translates high-level operations into optimized lower-level kernels for better performance of batch sampling (i.e., where the same computation is performed repeatedly). In addition, the package leverages JAX’s autovectorization (e.g., vmap) to automatically parallelize computations across candidate randomizations, further improving efficiency in large-scale designs. To further reduce memory use, computations can take advantage of low-precision data types, such as 16-bit floats for operations like matrix multiplication or aggregation operations, which can help conserve memory and accelerate numerical operations without significantly impacting the rerandomization criteria.
@@ 3.4 Hardware Acceleration
To further address the computational demands of rerandomization involving many units or covariates, fastrerandomize leverages GPU acceleration to mitigate runtime bottlenecks during both randomization generation and randomization-based inference. GPUs are particularly well-suited for parallelizing the small, repeated tasks that characterize rerandomization, such as repeatedly sampling and evaluating balance statistics for large pools of candidate randomizations.

### v2

# fastrerandomize: Fast Rerandomization Using Accelerated Computing

@@ Abstract
We introduce fastrerandomize, an R package that implements novel algorithmic approaches to rerandomization in experimental design. Rerandomization improves precision by discarding treatment assignments until covariate balance meets predefined thresholds, but existing implementations often struggle with computational demands in large-scale settings. fastrerandomize addresses these limitations through three key innovations: (1) optional hardware acceleration via GPU/TPU backends, (2) memory-efficient key-based storage that avoids explicit randomization storage, and (3) computational optimizations through auto-vectorization and just-in-time compilation. This algorithmic framework enables exact or Monte Carlo generation of rerandomized designs even with billions of candidate randomizations and stringent balance thresholds. Simulations demonstrate substantial performance gains over existing implementations (greater than 34-fold speedup reported here), particularly in high-dimensional settings. By integrating modern computational techniques with principled statistical methods, fastrerandomize extends the practical applicability of rerandomization to experimental designs that were previously computationally intractable. An interactive demonstration is available at fastrerandomize.github.io/space.
@@
Rerandomization, randomization tests, experimental design, covariate balance, hardware acceleration, computational methods
@@ Introduction
Rerandomization is a powerful strategy for ensuring good covariate balance in experimental research. By repeatedly drawing treatment assignments until a prespecified balance criterion is met, rerandomization helps improve the precision of estimated treatment effects relative to pure random assignment (Morgan and Rubin, 2012; Li et al., 2018). Despite its advantages, rerandomization poses a formidable computational challenge: as the size of an experiment grows, the space of possible assignments can quickly reach hundreds of billions, complicating both the generation of acceptable randomizations and the conduct of valid inference.
Although theoretical work has established the statistical validity of rerandomization, user-friendly software that efficiently implements these methods in large-scale settings has, until now, been more limited. Moreover, despite the statistical benefits of rerandomization in experimental design, existing implementations often struggle with computational bottlenecks as experiments scale. This is particularly true in the increasingly common high-dimensional RCT settings involving text, image, video, or network data.
The fastrerandomize package addresses these limitations by integrating modern computational techniques with these principled statistical methods. The package leverages three key innovations not (to our knowledge) seen in prior approaches: (1) hardware acceleration via GPU/TPU backends that delivers speedups greater than 34-fold over existing implementations, (2) memory-efficient key-based storage that reduces memory requirements in proportion to the sample size, and (3) advanced computational optimizations through auto-vectorization and just-in-time compilation. This algorithmic framework extends rerandomization to previously intractable scales, enabling designs even when assignment pools reach billions of combinations and stringent acceptance thresholds are desired. By bridging new computational methods with statistical theory, fastrerandomize makes rerandomization practically feasible for large-scale experimental research that was previously constrained by computational bottlenecks and resource limitations.
@@
In what follows, we provide an overview of the statistical framework underpinning fastrerandomize, detail its core functionalities, and illustrate how it may be applied in practice. After literature review in Section 1, Section 2 outlines the primary methodologies implemented in the package, Section 3 describes its software architecture and usage, with Section 4 performing a simulation study. Sections 5 and 6 offer examples demonstrating both basic and more advanced analyses. We conclude with potential directions for future development in Section 8. We provide an interactive web-based demonstration of the fastrerandomize package, accessible online at fastrerandomize.github.io/space, allowing users to explore rerandomization designs and tradeoffs in a web-only interface.
@@
Figure 1: Core workflow of the fastrerandomize package for conducting rerandomization and randomization testing. The process begins with covariate data and proceeds through five main steps: (1) generating candidate randomizations using compiled code, (2) checking covariate balance with GPU-accelerated computations, (3) accepting or rejecting assignments using Monte Carlo methods for large experiments, (4) conducting exact randomization tests using the pool of accepted randomizations, and (5) optionally computing fiducial intervals. Steps are implemented with compiler-level optimizations and hardware acceleration to enable efficient processing of large-scale experiments.
@@ 1 Literature Review and Related Packages
Over the past decade, an expanding body of open-source software has emerged to facilitate experimental design, with a particular focus on random assignment procedures and randomization-based inference. Many of these tools incorporate rerandomization techniques to improve covariate balance across treatment groups, thereby enhancing the statistical power of a study (Morgan and Rubin, 2012; Li et al., 2018). Despite the proliferation of solutions, existing packages often address only narrow classes of experiments and frequently encounter computational challenges in large-scale settings with many units or covariates. This section surveys some of the most widely used packages for general random assignment, randomization testing, and rerandomization, before discussing the core challenges that persist in these domains.
@@ 1.1 General Random Assignment and Testing Packages
A variety of open-source packages are available for general-purpose random assignment and randomization-based testing. The randomizr package, for instance, streamlines the assignment of study units to diverse treatment groups. Another solution, RATest, incorporates randomization tests applicable to multiple experimental design scenarios and offers sample datasets that illustrate, for example, permutation testing in regression discontinuity contexts (Olivares-Gonzalez and Sarmiento-Barbieri, 2017).
@@ 1.3 Challenges and Gaps
Although these packages advance the design and analysis of experiments, several limitations constrain their application to large-scale studies. One persistent concern is that many solutions can verify or test an existing randomization but provide limited functions for the initial task of constructing assignment schemes automatically. This requires researchers to generate and manage their own candidate pools of randomizations. If researchers desire stringent balance, many randomizations with poor balance need to be rejected along the way.
Scalability therefore presents a challenge in rerandomization: as experiments grow to include more units and covariates, generating, evaluating, and storing large pools of acceptable randomizations becomes computationally prohibitive without specialized algorithmic approaches. Current software packages often struggle when handling scenarios where investigators seek stringent acceptance thresholds with hundreds of billions of potential randomizations, or in experiments involving hundreds or thousands of covariates—a scenario becoming increasingly common with the rise of causal machine learning involving high-dimensional text (Keith et al., 2020), image (Jerzak et al., 2023), or network data (Ogburn et al., 2024).
Despite the availability of compute gains through compilation and autovectorization and parallel processing advantages in modern hardware, existing rerandomization implementations—particularly in the R ecosystem—have not taken advantage of these advances. The fastrerandomize package therefore addresses these limitations by automating balanced assignment generation and harnessing this acceleration, thereby making rerandomization and randomization-based inference both computationally efficient and practically feasible for large-scale experimental research.
@@ 2 Statistical Methods
Conversely, overly lenient thresholds can diminish the advantages of rerandomization by accepting many poorly balanced designs, causing the method to converge back to classical complete randomization. Accordingly, researchers often iterate over plausible acceptance probabilities in pre-analysis simulations to gauge how different thresholds affect downstream power, $p$-values, or interval widths (Branson et al., 2024; Kapelner et al., 2022). Such simulations also help ensure that the final pool of accepted randomizations is large enough for reliable test-based inferences and for constructing credible fiducial intervals, particularly in high-dimensional covariate settings (e.g., text, image, or network data (Keith et al., 2020; Jerzak et al., 2023; Ogburn et al., 2024)). [^1] In high-dimensional applications, one can improve numerical stability by incorporating penalized or regularized covariance estimates (e.g., diagonal or ridge-regularized inverses) into the balance statistic, ensuring the acceptance criterion continues to reflect meaningful imbalances.
@@ 3 Computational Methods
As can be seen from the above, implementing rerandomization at scale involves navigating an inherently combinatorial process, in which the total number of possible treatment assignments can be enormous.
To address these computational challenges, fastrerandomize employs a hybrid of exact enumeration (for smaller designs) and Monte Carlo sampling (for large designs) while leveraging batched processing for memory efficiency. Central to this approach is the use of key-based storage, which replaces the explicit storage of every randomization vector with small keys that allow the randomizations to be regenerated on demand. Finally, the package exploits just-in-time (JIT) compilation and hardware acceleration via GPUs or specialized backends to further reduce run times. The methods outlined here enable fastrerandomize to handle rerandomization tasks that would be computationally prohibitive using naive approaches, even when the number of units, covariates, or candidate randomizations is very large.
@@ 3.1 Exact Enumeration and Batched Processing
We first discuss how the package addresses a central challenge in rerandomization related to the combinatorial explosion of possible treatment assignments and associated memory constraints.
To tackle this challenge, the package provides two primary strategies for generating and storing randomizations—exact enumeration and Monte Carlo sampling—each designed to accommodate different experimental sizes and memory constraints. Both exploit a batched workflow to reduce memory overhead, while R-facing functions manage data validation, type conversion, and user interaction. This integrated approach enables researchers to efficiently construct pools of acceptable randomizations across a range of experimental scales.
@@
For smaller studies where the total number of possible assignments, $\binom{n}{n_{\textrm{Treated}}},$ is sufficiently small (e.g., up to a few hundred million), it is possible to exhaustively enumerate every distinct randomization. Users can invoke exact enumeration by specifying randomization_type=’’exact’’ in generate_randomizations(). This approach systematically iterates over all assignments, providing the maximum level of precision and guaranteeing that the entire randomization space is captured. Although exact enumeration is attractive from a statistical perspective, it becomes infeasible for even moderate sample sizes (e.g., $n>25$), because the number of possible assignments grows combinatorially.
@@
For larger experiments, fastrerandomize supports a Monte Carlo mechanism that draws candidate randomizations in batches, then evaluating their covariate balance. Under Monte Carlo sampling, candidate assignments are randomly generated and checked against a user-specified criterion (e.g., Hotelling’s $T^{2}$). Those randomizations meeting the threshold are accepted. This approach can rapidly explore massive assignment spaces without enumerating them in full and is typically favored for large $n$, and can be accessed setting randomization_type=’’monte_carlo’’ in generate_randomizations().
Memory Management.
Regardless of whether the randomizations are generated by exact enumeration or Monte Carlo sampling, fastrerandomize relies on batching to handle large-scale experiments efficiently. The package implements this batching strategy through two key mechanisms.
First, both the exhaustive listing of assignments and the Monte Carlo draws can be subdivided into user-defined batches, a process controlled by the batch_size parameter. This chunked generation approach prevents the software from loading all candidate randomizations simultaneously into memory, thereby mitigating potential memory constraints.
Second, the package offers flexible return options wherein users can further reduce memory loads by writing accepted assignments directly to disk rather than returning them to R, an operation facilitated by the file parameter in generate_randomizations(). This disk-writing capability proves especially valuable when generating hundreds of thousands or millions of accepted randomizations that would otherwise overwhelm available RAM.
By selecting exact enumeration for small designs and Monte Carlo with batching for large ones, analysts can scale rerandomization from modest experiments to those involving thousands or millions of units. This unified approach provides both the precision afforded by thorough enumeration and the computational tractability of sample-based generation, all while leveraging batched execution to prevent overwhelming system memory.
3.2 Key-Only Storage
Even with Monte Carlo sampling, storing all candidate randomizations in memory quickly becomes infeasible as $n$ and the total number of candidate randomizations grow. To address this, our implementation employs an approach where each batch of candidate randomizations is generated and balance evaluated, but only key information from the pseudo-random number generator (PRNG) generating the randomization vector is preserved for the final evaluation step. This keeps the memory footprint low. The result is a scalable procedure that supports large experiments, since we avoid loading or caching an unwieldy number of candidate assignments in a single data structure. This key-generation process is visualized in Figure 3.
@@ 3.2 Key-Only Storage
3.2.1 A Model of Memory Gains from Key-Only Storage
@@
Instead of storing every candidate randomization vector directly, one may store only an associated key $\kappa^{(m)}$ for each $W^{(m)}$. Here, $\kappa^{(m)}\in\{\mathbb{N}_{0}\}^{L}$ is a short seed (or split key) that can be used by a PRNG to regenerate the corresponding random vector $W^{(m)}$ on demand:
@@
For an experiment with 1000 units, this ratio would be $1000/2$, indicating memory usage improvements of 500 times. At any later point, to reconstruct the accepted $W^{(m)}$ values for final analysis, we simply apply the PRNG again to these stored keys. Memory gains as a function of sample size are visualized in Figure 4.
@@
Figure 4: Memory gains from key-only storage grow with sample size ($n$).
@@ 3.3 Hardware Acceleration
To further address the computational demands of rerandomization involving many units or covariates, fastrerandomize leverages GPU acceleration to mitigate runtime bottlenecks during both randomization generation and randomization-based inference. GPUs are particularly well-suited for parallelizing the small, repeated tasks that characterize rerandomization, such as repeatedly sampling and evaluating balance statistics for large pools of candidate randomizations. (For a model of rerandomization speedups from computations on GPU, see Appendix A).
@@
3.4 Just-in-Time Compilation with Precision Management
Finally, beyond memory-efficient storage techniques and optional hardware acceleration, fastrerandomize also accelerates computation through advanced compiler optimizations. Each internal package function is decorated with just-in-time (JIT) compilation, which translates high-level operations into optimized lower-level kernels for better performance of batch sampling (i.e., where the same computation is performed repeatedly). In addition, the package leverages autovectorization to automatically parallelize computations across candidate randomizations, further improving efficiency in large-scale designs. [^2] Autovectorization is a compiler optimization technique that automatically transforms sequential loops in code to use Single Instruction Multiple Data (SIMD) instructions, allowing multiple data elements to be processed simultaneously in parallel without requiring manual vectorization. To further reduce memory use, computations can also take advantage of low-precision data types, such as 16-bit floats for operations like matrix multiplication or aggregation operations, which can help conserve memory and accelerate numerical operations without significantly impacting the rerandomization criteria.
4 Simulation Study
4.1 Simulation Study Design
To assess the practical performance of the fastrerandomize package compared to base R implementations, we constructed a series of simulation studies that varied both sample size and the dimensionality of the covariate space. This analysis will also allow us to see the relative contribution of various settings within fastrerandomize (such as the optimal use of GPU acceleration).
In each replicate, we generate pre-treatment covariates from independent Gaussians, with sample size, $n\in\{10,100,1000\}$, and number of covariates, $k\in\{10,100,1000\}$. We randomize a binary treatment to half of units and simulate experimental outcomes under a linear model in which each unit’s potential outcomes depended on a random linear combination of the covariates plus additive Gaussian noise. We used Monte Carlo draws batched in blocks for memory efficiency. Ten replicates are used for each unique analysis arrangement.
For metrics, we focus on the computation time of randomization generation and randomization testing as the number of observations and covariates vary. We compare results across (a) a base R implementation, (b) the fastrerandomize implementation using a CPU backend, and (c) the fastrerandomize implementation using a GPU backend. For reasons that will become apparent, the base R implementation suffers in large $n$ and $k$ cases; as a result, we only report results for the smaller compute regimes.
4.2 Simulation Study Results
Figure 6 reports results for $n=100$. Even at relatively modest sample sizes, some patterns emerge. First, both CPU- and GPU-based fastrerandomize outperform the base R implementation, particularly as the covariate dimension grows. For instance, at $2\times 10^{5}$ Monte Carlo draws with 100 covariates, the fastrerandomize-GPU backend finishes in 5 seconds for randomization generation, whereas the base R approach exceeds 170 seconds (a 34-fold speedup for fastrerandomize). Moreover, fastrerandomize-GPU also outperforms the CPU-only-approach by up to 41% in the high-dimensional setting ($k=1000$). This result underscores the fact that even in small-sized experiments ($n=100$), compilation, streamlined memory management, and hardware acceleration substantially reduce overhead for scaling rerandomization and, therefore, improves iteration time for researchers.